OpenStudy (anonymous):

Can someone Explain step-by-step how to solve this? Find the equation of the perpendicular line to each of the following that passes through the point indicated. 4x + 5y = - 40 passing through the point (-1, 3)

OpenStudy (anonymous):

First put the first line into slope-intercept form. Do you know how to do that?

OpenStudy (anonymous):

isnt it y = MX + B?

OpenStudy (anonymous):

so it would be y = -4/5X - 8?

OpenStudy (anonymous):

are their any options?

OpenStudy (anonymous):

y = -4/5x + 8 soo yea your right!(:

OpenStudy (anonymous):

Options like answers? Yeah -5x + 4y = 17 5x + 4y = -17 5x - 4y = -17 5x + 4y = 17

OpenStudy (anonymous):

oh then so you have to turn it into standard form

OpenStudy (anonymous):

So how do i put it into Standard?

OpenStudy (anonymous):

Ax + By = C

OpenStudy (anonymous):

I got 4x + 5y = -40....I moved over the 4/5th to the left side then multiplied everything by 5 to get rid of the fraction...What did i do wrong? :/

OpenStudy (anonymous):

i really don't know i wish i could help you, try asking this question again and do you have flvs?

OpenStudy (anonymous):

florida virtual school? because this question looked familiar but im guessing i forgot how to do this.. i took flvs last year thats why

OpenStudy (anonymous):

oh haha well their both virtual schools. Good luck and sorry that i have been no help

OpenStudy (anonymous):

its cool thanks for trying

OpenStudy (anonymous):

Step 1: find the slope of given line. Step 2: think of relationship between two lines that is if two lines are parallel, the slopes are the same; if two lines are perpendicular then the slope of unknow line is negative reciprocal of given line. Step 3: use the slope you find in step 2 and the given point, substitute into point-slope form, \[y-y_1=m(x-x_1)\] step 4: write your equation in standard form (since the given line is in standard form)